# Cleaning WASP-33 b transits from the host star photometric variability: analysis of TESS data from two sectors

### v2

@@ 4.2 Transit lightcurves fit
$i$ [] $87.21(48)$ $\rho_{\star}$ [$\rho_{\odot}$] $0.4342(59)$ $A_{\rm I}$ $0.1883$
$r/R_{\star}$ $0.11078(25)$∗∗ $B_{\rm I}$ $0.2873$
$r$ [$R_{\rm Jup}$] $1.665(10)$ $A_{\rm II}$ $0.246(6)$
$a$ [AU] $0.02555(35)$ $B_{\rm II}$ $0.252(6)$
${}^{*\phantom{**}}$ Epoch $\rm BJD_{TDB}=2459000$.
${}^{**\phantom{*}}$ Taken from (Collier Cameron et al., 2010).
∗∗∗ This value is considerably affected by the third light correction, so its uncertainty may be unreliable.
Fitting the detrended and whitened ITD allowed us to obtain an improved set of WASP-33 system parameters. They are given in Table 3. This fit used the stellar mass estimate by Collier Cameron et al. (2010), but ignored their stellar radius value since we need to know the correlation between $M_{\star}$ and $R_{\star}$ to correctly include them both. Our estimate for the stellar density appeared somewhat smaller though (by $\sim 2.5\sigma$). This model assumes that transit timings strictly follow a linear ephemeris (no TTV).
In addition to the best fitting estimates of limb darkening coefficients, we also give in Table 3 their reference values (i) computed from the tables by Claret (2017) using $T_{\rm eff}$, $\log g$, and [Fe/H] by Collier Cameron et al. (2010) and (ii) computed by von Essen et al. (2020). These reference sets of estimates both differ from the fitted one, as well as from each other. This may indicate some issues regarding their reliability for not yet clear reasons.

### v3

@@ 4.2 Transit lightcurves fit
$i$ [] $87.21(48)$ $\rho_{\star}$ [$\rho_{\odot}$] $0.4342(59)$ $A^{\rm C17}$ $0.1883$∗∗
$r/R_{\star}$ $0.11078(25)$∗∗∗ $M_{\star}^{\rm DR3}$ [$M_{\odot}$] $1.54(4)$∗∗ $B^{\rm C17}$ $0.2873$∗∗
$r$ [$R_{\rm Jup}$] $1.665(10)$ $R_{\star}^{\rm DR3}$ [$R_{\odot}$] $1.59(3)$∗∗ $A^{\rm vE20}$ $0.246(6)$∗∗
$a$ [AU] $0.02555(35)$ $B^{\rm vE20}$ $0.252(6)$∗∗
${}^{*\phantom{***}}$ Epoch $\rm BJD_{TDB}=2459000$.
${}^{**\phantom{**}}$ Taken from (Collier Cameron et al., 2010).
${}^{***\phantom{*}}$ Not used in the fit and given for comparison only (see text).
∗∗∗∗ This value is considerably affected by the third light correction, so its uncertainty may be unreliable.
Fitting the detrended and whitened ITD allowed us to obtain an improved set of WASP-33 system parameters, given in Table 3. This model assumes that transit timings strictly follow a linear ephemeris (no TTV). The fit included the stellar mass estimate by Collier Cameron et al. (2010) which was based on their spectrum model, line-profile tomography, and ground-based transit data constraining stellar density. This stellar mass was accounted for through a penalty function added to the log-likelihood (see Galazutdinov et al., 2023). We ignored the accompanying stellar radius value since we need to know the correlation between $M_{\star}$ and $R_{\star}$ to correctly include them both. Our estimate for the stellar density appeared smaller by $\sim 2.5\sigma$ than that by Collier Cameron et al. (2010).
Independent estimations of WASP-33 mass and radius are available in GAIA DR3 FLAME data table (Gaia Collaboration; Vallenari et al., 2023; Gaia Collaboration, 2024); they were added to Table 3 for comparison. Although these estimations use more accurate parallax, we did not use them in our fit, because contrary to Collier Cameron et al. (2010) they (i) are based on a rather low-resolution spectrum, (ii) did not include constraints from any transit observations of WASP-33 b. This resulted in a considerably larger uncertainties in GAIA $M_{\star}$, $R_{\star}$. We were also bothered by very low metallicity of $-0.56$ given in the same GAIA table and looking inconsistent with e.g. Collier Cameron et al. (2010). Yet another independent $M_{\star}$, $R_{\star}$ estimations provided by von Essen et al. (2020) have even larger uncertainties than GAIA ones.
In addition to the best fitting estimates of limb darkening coefficients, we also give in Table 3 their reference values (C17) computed from the tables by Claret (2017) using $T_{\rm eff}$, $\log g$, and [Fe/H] by Collier Cameron et al. (2010) and (vE20) computed by von Essen et al. (2020). These reference sets of estimates both differ from the fitted one, as well as from each other. This may indicate some issues regarding their reliability for not yet clear reasons.